OpenStudy (anonymous):

solve for x: (log[3](log[3]x))=-3

hartnn (hartnn):

\(\huge if \quad log_ax=b \quad then \quad a^b=x\) so what will be (log[3](log[3]x))=-3 ?

hartnn (hartnn):

so u get (-3)^3 = ?

OpenStudy (anonymous):

x=-3^3 =>-27? i still don't understand what happens to the other log[3]x?

hartnn (hartnn):

(log[3](log[3]x))=-3 gives u log[3]x=(-3)^3 using definition of log i mentioned in 1st comment. got this.?

OpenStudy (anonymous):

oh okay I got it.

hartnn (hartnn):

so whats your final answer ?